40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Feb 7 15:00:16 2020, gautam, Update, LSC, More high BW POY experiments AO_TFs.pdf
    Reply  Fri Feb 7 19:39:10 2020, Koji, Update, LSC, More high BW POY experiments 
Message ID: 15199     Entry time: Fri Feb 7 15:00:16 2020     Reply to this: 15200
Author: gautam 
Type: Update 
Category: LSC 
Subject: More high BW POY experiments 

To study the evilution of the AO path TFs a bit more, I've hooked up POY11_Q Mon to IN1 of the CM board. I will revert the usual setup later in the evening.

Update 1730: I've returned the cabling at 1Y2 to the nominal config, and also reverted all EPICS settings that I modified for this test. Y-arm POY locking works. Attachment #1 shows the summary of the results of this test - note that the AO gain was kept fixed at +5dB throughout the test. I have arbitrarily trimmed the length of the frequency vector for some of these traces so that the noisy measurement doesn't impede visual interpretation of the plots so much. At first glance, the performance is as advertised. I basically followed the settings I had here to get started, and then ramped up various gains to check if the measured OLTF evolved in the way that I expected it to. The phase lead due to the AO path is clearly visible.

Some important differences between this test and the REFL11 blending is (i) in the latter case, there will also be a parallel loop, CARM_A, which is effecting some control, and (ii) the optical gain of CARM-->REFL11_I is much higher than L_Y-->POY. So the initial gain settings will have to be different. But I hope to get some insight into what the correct settings should be from this test. I think IMC servo IN2 gain and AO gain slider on the CM board are degenerate in the effect they have, modulo subtle effects like saturation.

One possibility is that the gain allocation I used yesterday was wrong for the dynamic range of the CARM error signal. In some initial trials today, when I set the CM board IN1 gain to -32dB (as in the case of attempting the CARM RF handoff) and compensated for the reduced POY PDH fringe amplitude by increasing the digital gain for the CM_Slow path, I found that there was no phase advance visible even when I ramped up the IMC IN2 gain to +10dB. So, for the CARM handoff too, I might have to start with a higher CM board IN_1 gain, compensate by reducing the CM_Slow digital gain even more, and then try upping the IMC IN2 gain.

P.S. When the excitation input to the CM board was enabled in order to make TF measurements, I saw significant increase in the RMS of the error signal. Probably some kind of ground loop issue.

Attachment 1: AO_TFs.pdf  244 kB  Uploaded Fri Feb 7 18:57:46 2020  | Hide | Hide all
AO_TFs.pdf
ELOG V3.1.3-